OpenStudy (anonymous):

!!! Help Please Determine if the function f(x)=x^2l5xl is odd, even, or neither. I need help to start and finish it. please help

jimthompson5910 (jim_thompson5910):

are you able to find f(-x) ?

OpenStudy (anonymous):

Do you know the first step?

OpenStudy (anonymous):

My notes are just not of help. and f(-x) do I just replace the x with a -x?

jimthompson5910 (jim_thompson5910):

every x with -x, yes

OpenStudy (anonymous):

Im sure I did this wrong but I think -5x^3?

jimthompson5910 (jim_thompson5910):

f(x)=x^2l5xl f(-x)=(-x)^2l5(-x)l f(-x)=(-x)^2l-1*5xl f(-x)=(-x)^2l-1|*|5xl f(-x)=(-x)^2*(1)*|5xl f(-x)=x^2*|5xl

jimthompson5910 (jim_thompson5910):

So because f(-x) = f(x), we can say that f(x) is an even function

OpenStudy (anonymous):

Ohh ok I got it and took note thank you!